The demand function for a particular good in competitive market is given by y = a + bP. What are the associated direct and indirect utility functions?


1
Expert's answer
2021-12-13T10:23:28-0500

For the given function of demand  "y=a+bp"

 The function for "indirect utility" can be written as 

"V(y,p,m)=max u(y,z)"

where, 

z=other goods

m=income/wealth

Indirect utility function"=−\frac{\frac{∂V}{∂p}}{\frac{∂V}{∂m}}"

Indirect utility function"=y(p,m)"


Thus, the consumer can only get as much utility as limited to the price and also to their income/wealth. 

Now, for the utility function of "direct" nature, 

"U(y,z)=Min V(y,p,m)"